OpenStudy (anonymous):

Bridget has a kitchen canister, in the shape of a cylinder, which holds flour. If the canister is 20 centimeters wide and 28 centimeters tall, what is its approximate surface area in terms of pi? Answer 1,920π square centimeters 1,120π square centimeters 760π square centimeters 560π square centimeters

sam (.sam.):

suface area of a cylinder = 2πrh where r = radius, d/2, width ?/2 h = height A = 2π(20/2)(28) A = 560π square centimeters

OpenStudy (anonymous):

so this one is like where you use \[2\pi rh + 2\pi r ^{2}\] ??

OpenStudy (anonymous):

i think what you put would be the lateral area

sam (.sam.):

Ah, yeah use your formula 2πrh+2πr^2 A = 560π +200π A = 760π

sam (.sam.):

C:) 760π square centimeters

sam (.sam.):